# Sum of Squares Decompositions and Rank Bounds for Biquadratic Forms

### v2

@@ Abstract
We study positive semi-definite (PSD) biquadratic forms and their sum-of-squares (SOS) representations. For the class of partially symmetric biquadratic forms, we establish necessary and sufficient conditions for positive semi-definiteness and prove that every PSD partially symmetric biquadratic form is a sum of squares of bilinear forms. This extends the known result for fully symmetric biquadratic forms. We describe an efficient computational procedure for constructing SOS decompositions, exploiting the Kronecker-product structure of the associated matrix representation. We present a $2\times 2$ PSD biquadratic form and show that it can be expressed as the sum of three squares, but cannot be expressed as the sum of two squares. Furthermore, we present a $3\times 2$ PSD biquadratic form and show that it can be expressed as the sum of four squares, but cannot be expressed as the sum of three squares. This shows that previously proved results that a $2\times 2$ PSD biquadratic form can be expressed as the sum of three squares, and a $3\times 2$ PSD biquadratic form can be expressed as the sum of four squares, are tight. Moreover, we establish a universal upper bound $\operatorname{SOS\text{-}rank}(P)\leq mn-1$ for any SOS biquadratic form, which improves the trivial bound $mn$ and is tight in small dimensions.
Keywords. Biquadratic forms, sum-of-squares, positive semi-definiteness, M-eigenvalues, partially symmetric biquadratic forms.
@@ 1 Introduction
In [5], it was proved that a $2\times 2$ positive semi-definite biquadratic form can be expressed as the sum of three squares. In [13], it was proved that a $3\times 2$ positive semi-definite biquadratic form can be expressed as the sum of four squares. In this paper, we show that these two results are tight by presenting a $2\times 2$ positive semi-definite biquadratic form which cannot be expressed as the sum of two squares, and a $3\times 2$ positive semi-definite biquadratic form which cannot be expressed as the sum of three squares. These results, to the best of our knowledge, are the first two SOS rank tightness results for biquadratic forms.
@@ 4 Extension to General Partially Symmetric Forms
5 Tightness of the SOS Rank Bounds for $2\times 2$ and $3\times 2$ Forms
@@
Suppose that $P$ admits an SOS decomposition $\sum_{r=1}^{R}f_{r}^{2}$. Then, it holds that $\text{supp}(f_{r})\subseteq\frac{1}{2}\text{New}(P)$ [15, Theorem 1][5, Theorem 4]. Here, $\text{New}(P)$ is the Newton polytope of $P$. From this we may derive that the SOS decomposition can only be the SOS of bilinear forms, i.e., $\sum_{r}({\bf c}_{r}^{\top}{\bf z})^{2}$. Thus, the SOS rank of $P$ equals the minimum possible rank of $M(\Gamma)$ satisfying $M(\Gamma)\succeq 0$. ∎
Theorem 1 of [5] shows that a $2\times 2$ PSD biquadratic form can always be expressed as the sum of three squares. The following theorem shows that this bound is tight.
Theorem 5.2 (Tightness of the bound for $2\times 2$ forms).
The biquadratic form
$P(\mathbf{x},\mathbf{y})=3x_{1}^{2}y_{1}^{2}+2x_{1}^{2}y_{2}^{2}+2x_{2}^{2}y_{1}^{2}+3x_{2}^{2}y_{2}^{2}$
is positive semidefinite and satisfies $\operatorname{SOS-rank}(P)=3$. An explicit sum-of-squares decomposition is
$P(\mathbf{x},\mathbf{y})=$ $\left(\sqrt{3}\,x_{1}y_{1}-\frac{2\sqrt{3}}{3}\,x_{2}y_{2}\right)^{2}+\left(\frac{\sqrt{15}}{3}\,x_{2}y_{2}\right)^{2}$
$+$ $\left(\sqrt{2}\,x_{1}y_{2}+\sqrt{2}\,x_{2}y_{1}\right)^{2}.$
Moreover, $P$ cannot be written as a sum of two squares of bilinear forms, showing that the upper bound $\operatorname{SOS-rank}(P)\leq 3$ for $2\times 2$ PSD forms is tight.
Proof.
One can verify by direct computation that the given SOS decomposition holds. We now prove that no SOS decomposition can consist of only two squares. Consider the symmetric matrices
$A=\begin{pmatrix}3&0&0&-1\\ 0&2&1&0\\ 0&1&2&0\\ -1&0&0&3\end{pmatrix},\hskip 28.80008ptD=\begin{pmatrix}0&0&0&1\\ 0&0&-1&0\\ 0&-1&0&0\\ 1&0&0&0\end{pmatrix}$.
Define ${\bf z}=(x_{1}y_{1},x_{1}y_{2},x_{2}y_{1},x_{2}y_{2})^{\top}$. A direct computation from (12) shows that $P$ can be expressed by the quadratic form in $\bf z$} with exactly one free parameter $\lambda\in\mathbb{R}$, i.e.,
$P({\bf x},{\bf y})={\bf z}^{\top}\bigl(A+\lambda D\bigr){\bf z}.$
Thus, by (13), the SOS rank of $P$ equals the minimum possible rank of $A+\lambda D$ as $\lambda$ varies over $\mathbb{R}$ and $A+\lambda D\succeq 0$; i.e.,
$\operatorname{SOS\text{-}rank}(P)=\min_{\lambda\in\mathbb{R},A+\lambda D\succeq 0}\;\operatorname{rank}(A+\lambda D).$
To compute this minimum, observe that the orthogonal matrix
$B=\begin{pmatrix}0&-\frac{\sqrt{2}}{2}&\frac{\sqrt{2}}{2}&0\\[2.0pt] \frac{\sqrt{2}}{2}&0&0&\frac{\sqrt{2}}{2}\\[2.0pt] 0&\frac{\sqrt{2}}{2}&\frac{\sqrt{2}}{2}&0\\[2.0pt] -\frac{\sqrt{2}}{2}&0&0&\frac{\sqrt{2}}{2}\end{pmatrix}$
satisfies
$BAB^{\top}=\operatorname{diag}(1,2,3,4),\hskip 28.80008ptBDB^{\top}=\operatorname{diag}(1,1,-1,-1).$
Hence $A+\lambda D$ is congruent to
$\operatorname{diag}(1+\lambda,\;2+\lambda,\;3-\lambda,\;4-\lambda).$
For any real $\lambda$, at most one diagonal entry of this matrix can be zero. Thus, $\operatorname{rank}(A+\lambda D)\geq 3$ for all $\lambda\in\mathbb{R}$, and the minimum is attained at, for example, $\lambda=-1$, yielding $\operatorname{rank}(A-D)=3$.
The explicit decomposition into three squares shows that this lower bound is attainable, completing the proof. ∎
Similarly, Theorem 2 of [13] shows that a $3\times 2$ PSD biquadratic form can always be expressed as the sum of four squares. The following theorem shows that this bound is tight.
Theorem 5.3 (Tightness of the bound for $3\times 2$ forms).
The biquadratic form
$P(\mathbf{x},\mathbf{y})=\|\mathbf{x}\|^{2}\|\mathbf{y}\|^{2}=\sum_{i=1}^{3}x_{i}^{2}\sum_{j=1}^{2}y_{j}^{2}$
is positive semidefinite and satisfies $\operatorname{SOS-rank}(P)=4$. An explicit sum-of-squares decomposition is
$P(\mathbf{x},\mathbf{y})$ $=(x_{1}y_{1}+x_{2}y_{2})^{2}+(x_{1}y_{2}-x_{2}y_{1})^{2}$
$\quad+(x_{3}y_{1})^{2}+(x_{3}y_{2})^{2}.$
Moreover, $P$ cannot be written as a sum of three squares of bilinear forms, showing that the upper bound $\operatorname{SOS-rank}(P)\leq 4$ for $3\times 2$ PSD forms is tight.
Proof.
The validity of the provided SOS decomposition can be checked directly. We now proceed to demonstrate that no decomposition into three squares exists. Let $\mathbf{z}=(x_{1}y_{1},x_{1}y_{2},x_{2}y_{1},x_{2}y_{2},x_{3}y_{1},x_{3}y_{2})^{\top}\in\mathbb{R}^{6}$. Then $P(\mathbf{x},\mathbf{y})=\mathbf{z}^{\top}I_{6}\mathbf{z}=\|\mathbf{z}\|^{2}$.
A direct computation from (12) shows that $P$ can be expressed by the quadratic form in $\bf z$} with exactly three free parameters $a,b,c\in\mathbb{R}$, i.e.,
$P({\bf x},{\bf y})={\bf z}^{\top}(I_{6}+aA_{1}+bA_{2}+cA_{3}){\bf z},$
where
$A_{1}=\begin{pmatrix}0&0&0&1&0&0\\ 0&0&-1&0&0&0\\ 0&-1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{pmatrix},\qquad A_{2}=\begin{pmatrix}0&0&0&0&0&1\\ 0&0&0&0&-1&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&-1&0&0&0&0\\ 1&0&0&0&0&0\end{pmatrix}$,
$A_{3}=\begin{pmatrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&-1&0\\ 0&0&0&-1&0&0\\ 0&0&1&0&0&0\end{pmatrix}$.
By (13), it holds that
$\operatorname{SOS-rank}(P)=\min_{\begin{subarray}{c}a,b,c\in\mathbb{R}\\ I_{6}+S\succeq 0\end{subarray}}\operatorname{rank}\bigl(I_{6}+S\bigr).$
where $S=aA_{1}+bA_{2}+cA_{3}$.
The characteristic polynomial of $S$ is
$\det(\lambda I_{6}-S)=\lambda^{2}\bigl(\lambda^{2}-(a^{2}+b^{2}+c^{2})\bigr)^{2}.$
Thus the eigenvalues of $S$ are 0 (multiplicity 2) and $\pm\sqrt{a^{2}+b^{2}+c^{2}}$ (each with multiplicity 2). Consequently, the eigenvalues of $I_{6}+S$ are
[…]
each of the last two with multiplicity 2.
For $I_{6}+S$ to be positive semidefinite we need $1-\sqrt{a^{2}+b^{2}+c^{2}}\geq 0$, i.e., $a^{2}+b^{2}+c^{2}\leq 1$. Under this condition, the only possible zero eigenvalue occurs when $\sqrt{a^{2}+b^{2}+c^{2}}=1$. In that case the eigenvalue $1-\sqrt{a^{2}+b^{2}+c^{2}}$ becomes zero, giving two zero eigenvalues (multiplicity 2). Therefore
$\min_{\begin{subarray}{c}a,b,c\in\mathbb{R}\\ I_{6}+aA_{1}+bA_{2}+cA_{3}\succeq 0\end{subarray}}\operatorname{rank}\bigl(I_{6}+aA_{1}+bA_{2}+cA_{3}\bigr)=4.$
Thus $\operatorname{SOS-rank}(P)=4$. Since the upper bound for $3\times 2$ PSD biquadratic forms is also 4 [13], this shows the bound is tight. ∎
6 An Improved Universal SOS Rank Bound for General Biquadratic Forms
@@
This bound is sharp for many small cases: for $2\times 2$ forms the maximal SOS rank is 3=413=4-1 (Theorem 5.2). However, for $3\times 2$ forms the maximal SOS rank is 4=624=6-2 (Theorem 5.3). Whether the bound mn1mn-1 is always attainable for $m,n\geq 3$ remains an open question.
Remark 6.3.
@@
$4\;\leq\;\operatorname{BRS}(3,3)\;\leq\;8.$
@@
Determining the exact value of $\operatorname{BRS}(3,3)$ whether it equals $4$, 88, or some intermediate integer remains an open problem.

### v3

@@ Abstract
We study positive semi-definite (PSD) biquadratic forms and their sum-of-squares (SOS) representations. For the class of partially symmetric biquadratic forms, we establish necessary and sufficient conditions for positive semi-definiteness and prove that every PSD partially symmetric biquadratic form is a sum of squares of bilinear forms. This extends the known result for fully symmetric biquadratic forms. We describe an efficient computational procedure for constructing SOS decompositions, exploiting the Kronecker-product structure of the associated matrix representation. We introduce simple biquadratic forms. For $m\geq 2$, we present a $m\times 2$ PSD biquadratic form and show that it can be expressed as the sum of $m+1$ squares, but cannot be expressed as the sum of $m$ squares. This provides a lower bound for sos rank of $m\times 2$ biquadratic forms, and shows that previously proved results that a $2\times 2$ PSD biquadratic form can be expressed as the sum of three squares, and a $3\times 2$ PSD biquadratic form can be expressed as the sum of four squares, are tight. We also present an $3\times 3$ SOS biquadratic form, which can be expressed as the sum of six squares, but not the sum of five squares. Moreover, we establish a universal upper bound $\operatorname{SOS\text{-}rank}(P)\leq mn-1$ for any SOS biquadratic form, which improves the trivial bound $mn$.
Keywords. Biquadratic forms, sum-of-squares, positive semi-definiteness, sos rank, M-eigenvalues, partially symmetric biquadratic forms.
@@ 1 Introduction
We introduce simple biquadratic forms. For $m\geq 2$, we present a $m\times 2$ PSD biquadratic form and show that it can be expressed as the sum of $m+1$ squares, but cannot be expressed as the sum of $m$ squares. This provides a lower bound for sos rank of $m\times 2$ biquadratic forms, and shows that previously proved results that a $2\times 2$ PSD biquadratic form can be expressed as the sum of three squares, and a $3\times 2$ PSD biquadratic form can be expressed as the sum of four squares, are tight. We also present an $3\times 3$ SOS biquadratic form, which can be expressed as the sum of six squares, but not the sum of five squares.
@@ 4 Extension to General Partially Symmetric Forms
5 Lower Bounds for SOS Rank in Small Dimensions
In this section, we use simple biquadratic forms as a tool to present some lower bounds for SOS rank in small dimensions.
5.1 Simple Biquadratic Forms
Let $m\geq n$. We say a biquadratic form is a simple biquadratic forms if it contains only distinct terms of the type $x_{i}^{2}y_{j}^{2}$. Then we define a simple biquadratic form series $P_{m,n,s}$ as follows. For each $s=1,2,\dots,mn$, the form $P_{m,n,s}$ contains exactly $s$ distinct terms of the type $x_{i}^{2}y_{j}^{2}$.
Assume that the terms are taken in the following order. For $k=0,1,\dots,mn-1$, write $k=pm+q$ with integers $0\leq p<n$ and $0\leq q<m$ (so $p=\lfloor k/m\rfloor$ and $q=k\bmod m$). Define
$i_{k}=q+1,\hskip 28.80008ptj_{k}={\big(p+q\big)}\bmod n+1.$
Then the sequence of index pairs is $(i_{0},j_{0}),(i_{1},j_{1}),\dots,(i_{mn-1},j_{mn-1})$. We set
$P_{m,n,s}({\bf x},{\bf y})=\sum_{k=0}^{s-1}x_{i_{k}}^{2}y_{j_{k}}^{2}.$
Thus $P_{m,n,s}$ has exactly $s$ distinct square terms, and $P_{m,n,mn}$ contains every term $x_{i}^{2}y_{j}^{2}$ ($i=1,\dots,m$, $j=1,\dots,n$) exactly once.
For the small dimensions used in our theorems, the resulting forms are:
Case $m=2,n=2$:
$P_{2,2,1}$ $=x_{1}^{2}y_{1}^{2},$
$P_{2,2,2}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2},$
$P_{2,2,3}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{1}^{2}y_{2}^{2},$
$P_{2,2,4}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}.$
Case $m=3,n=2$:
$P_{3,2,1}$ $=x_{1}^{2}y_{1}^{2},$
$P_{3,2,2}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2},$
$P_{3,2,3}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{1}^{2},$
$P_{3,2,4}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2},$
$P_{3,2,5}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2},$
$P_{3,2,6}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{3}^{2}y_{2}^{2}.$
Case $m=3,n=3$:
$P_{3,3,1}$ $=x_{1}^{2}y_{1}^{2},$
$P_{3,3,2}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2},$
$P_{3,3,3}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2},$
$P_{3,3,4}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2},$
$P_{3,3,5}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2},$
$P_{3,3,6}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}.$
5.2 Tightness Examples and a General Lower Bound
Theorem 1 of [5] shows that a $2\times 2$ PSD biquadratic form can always be expressed as the sum of three squares. Theorem 2 of [14] shows that a $3\times 2$ PSD biquadratic form can always be expressed as the sum of four squares. The following theorem shows that these two bounds are tight, and present a general lower bound for $m\geq n=2$.
Theorem 5.1 (A general lower bound for $m\times 2$ forms).
Let $m\geq 2$. The simple biquadratic form
$P_{m,2,m+1}({\bf x},{\bf y})=\sum_{k=0}^{m}x_{i_{k}}^{2}y_{j_{k}}^{2},$
where $(i_{k},j_{k})$ are defined by the ordering rule in Section 2, is positive semidefinite and satisfies
$\operatorname{SOS\text{-}rank}(P_{m,2,m+1})=m+1.$
In particular, $P_{m,2,m+1}$ cannot be written as a sum of $m$ squares of bilinear forms.
Proof.
The nonnegativity and the upper bound $\operatorname{SOS\text{-}rank}(P_{m,2,m+1})\leq m+1$ are immediate from the decomposition
$P_{m,2,m+1}=\sum_{k=0}^{m}(x_{i_{k}}y_{j_{k}})^{2}.$
For the lower bound, assume for contradiction that $P_{m,2,m+1}=\sum_{t=1}^{m}L_{t}^{2}$ with each $L_{t}$ bilinear. Write $L_{t}=\sum_{i=1}^{m}\sum_{j=1}^{2}c_{ij}^{(t)}x_{i}y_{j}$. Let $A_{i}=(c_{i1}^{(t)})_{t=1}^{m}\in\mathbb{R}^{m}$ and $B_{i}=(c_{i2}^{(t)})_{t=1}^{m}\in\mathbb{R}^{m}$.
From the square terms present in $P_{m,2,m+1}$ one obtains the following conditions:
  • For the index $i_{0}=1$, both $x_{1}^{2}y_{1}^{2}$ and $x_{1}^{2}y_{2}^{2}$ appear; hence $\|A_{1}\|^{2}=1$ and $\|B_{1}\|^{2}=1$.
  • For each $i=2,\dots,m$, exactly one of $x_{i}^{2}y_{1}^{2}$ or $x_{i}^{2}y_{2}^{2}$ appears, depending on the parity of $i$ in the ordering. Consequently, for each such $i$, either $\|A_{i}\|^{2}=1$ and $B_{i}=0$, or $\|B_{i}\|^{2}=1$ and $A_{i}=0$.
The form $P_{m,2,m+1}$ contains no mixed terms $x_{p}x_{q}y_{r}y_{s}$ with $p\neq q$ or $r\neq s$. Setting the coefficients of all such mixed monomials in $\sum_{t}L_{t}^{2}$ to zero yields the orthogonality relations
$A_{p}\cdot A_{q}=0\;(p\neq q),\hskip 28.80008ptB_{p}\cdot B_{q}=0\;(p\neq q),\hskip 28.80008ptA_{p}\cdot B_{q}=0\;(\text{all }p,q).$
Thus the nonzero vectors among $\{A_{1},B_{1},A_{2},B_{2},\dots,A_{m},B_{m}\}$ are pairwise orthogonal. Because of the square?term conditions, exactly $m+1$ of these vectors are nonzero and have norm $1$ (namely $A_{1}$, $B_{1}$, and one vector for each $i=2,\dots,m$). We therefore have $m+1$ mutually orthogonal unit vectors in $\mathbb{R}^{m}$, which is impossible. Hence no $m$-square decomposition exists, and $\operatorname{SOS\text{-}rank}(P_{m,2,m+1})\geq m+1$. ∎
Remark 5.2.
Fang and Huang [7] informed us that they obtained the same lower bound as the above theorem but with a different method.
5.3 A Lower Bound for SOS Rank of $3\times 3$ Biquadratic Forms
Denote the SOS rank of a bilinear form by $\operatorname{sos}(P)$.
Theorem 5.3 (A $3\times 3$ form requiring six squares).
Let
$P^{\prime}({\bf x},{\bf y})=P_{3,3,6}({\bf x},{\bf y})=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2},$
where ${\bf x}=(x_{1},x_{2},x_{3}),{\bf y}=(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}$. Then
$\operatorname{sos}(P^{\prime})=6.$
Proof.
The upper bound $\operatorname{sos}(P^{\prime})\leq 6$ is immediate by taking
$L_{1}=x_{1}y_{1},\;L_{2}=x_{2}y_{2},\;L_{3}=x_{3}y_{3},\;L_{4}=x_{1}y_{2},\;L_{5}=x_{2}y_{3},\;L_{6}=x_{3}y_{1},$
which yields $P^{\prime}=\sum_{i=1}^{6}L_{i}^{2}$.
For the lower bound, assume for contradiction that $P^{\prime}=\sum_{i=1}^{5}L_{i}^{2}$ with each $L_{i}$ bilinear. Write $L_{i}=\sum_{a,b=1}^{3}c_{ab}^{(i)}x_{a}y_{b}$. Because $P^{\prime}$ contains exactly the six square terms $x_{a}^{2}y_{b}^{2}$ for
$(a,b)\in S=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1)\}$ with coefficient $1$ and no other square terms, we must have
$\sum_{i=1}^{5}(c_{ab}^{(i)})^{2}=1\qquad((a,b)\in S),\hskip 28.80008pt\sum_{i=1}^{5}(c_{ab}^{(i)})^{2}=0\qquad((a,b)\notin S).$
Hence each $L_{i}$ involves only the six coefficient types corresponding to $S$:
$L_{i}=a_{i}x_{1}y_{1}+b_{i}x_{1}y_{2}+c_{i}x_{2}y_{2}+d_{i}x_{2}y_{3}+e_{i}x_{3}y_{3}+f_{i}x_{3}y_{1}.$
Define vectors $A,B,C,D,E,F\in\mathbb{R}^{5}$ by $A=(a_{1},\dots,a_{5})$, etc. From the square-term conditions we have $\|A\|=\|B\|=\cdots=\|F\|=1$.
Since $P^{\prime}$ contains no mixed terms $x_{p}x_{q}y_{r}y_{s}$ with $p\neq q$ or $r\neq s$, all cross terms in $\sum_{i=1}^{5}L_{i}^{2}$ must vanish. This yields the orthogonality conditions
$A\!\cdot\!B=A\!\cdot\!C=\dots=E\!\cdot\!F=0,$
i.e., the six vectors $A,\dots,F$ in $\mathbb{R}^{5}$ are pairwise orthogonal and each has norm $1$. This is impossible because $\mathbb{R}^{5}$ can contain at most five nonzero mutually orthogonal vectors. Therefore no such decomposition with five squares exists, and $\operatorname{sos}(P^{\prime})\geq 6$. ∎
6 An Improved Universal SOS Rank Bound for General Biquadratic Forms
@@
Suppose that $P$ admits an SOS decomposition $\sum_{r=1}^{R}f_{r}^{2}$. Then, it holds that $\text{supp}(f_{r})\subseteq\frac{1}{2}\text{New}(P)$ [16, Theorem 1][5, Theorem 4]. Here, $\text{New}(P)$ is the Newton polytope of $P$. From this we may derive that the SOS decomposition can only be the SOS of bilinear forms, i.e., $\sum_{r}({\bf c}_{r}^{\top}{\bf z})^{2}$. Thus, the SOS rank of $P$ equals the minimum possible rank of $M(\Gamma)$ satisfying $M(\Gamma)\succeq 0$. ∎
@@
$6\;\leq\;\mathrm{BSR}(3,3)\;\leq\;8,$
where the lower bound follows from Theorem 5.3.
@@
Determining the exact value of $\operatorname{BRS}(3,3)$ whether it equals $6$, 77 or $8$, remains an open problem.